What is the volume of 3.0 g of lead? The density o
Answer #1

4) we knwo that

Density = Mass / Volume

So volume = Mass / Density = 3 / 11.34 = 0.264 mL = 0.26 m L

5) we know that

1miles / hour = 0.447 m /s

So 850 miles / hr = 0.447 X 850 m /s = 379.95 m /s = 380 m/s

6) It is a chemical change as one compound is converted to two compounds

Water is not ionic

It is not a nuclear reaction

Not a physical change